OpenStudy (henryarias14):

can i say that (1-9x^2)^-1/2 = sqr(1-9x^2)?

OpenStudy (henryarias14):

\[(1-9x^2)^{-1/2} = \sqrt{(1-9x^2)}\]

OpenStudy (henryarias14):

or does the 1/2 has to be positive?

OpenStudy (henryarias14):

@sunnnystrong

OpenStudy (sunnnystrong):

Hmmm.. it has to be positive because: \[(1-9x^2)^{-1/2}=\frac{ 1 }{ \sqrt{1-9x^2} }\]